OpenStudy (anonymous):

solve the equation by completing the square. Round to the nearest hundredth if necessary. x^2 + 10x = 18 (1 point) –11.56, 1.56 11.56, 1.56 –11.56, –1.56 11.56, –1.56

OpenStudy (campbell_st):

add (b/2)^2 to both sides (10/2)^2 = 25 x^2 + 10x + 25 = 43 (x + 5)^2 = 43 then taking thesquare root of both sides \[x + 5 = \pm \sqrt{43}\] then \[x = -5 \pm \sqrt{43}\] you need to evaluate the last part correct to 2 decimal places.